OpenStudy (anonymous):

Solve, using the substitution method 4x + 2y = 11 x - 2 = -2y

OpenStudy (anonymous):

So far i got

OpenStudy (anonymous):

4(2 - 2y) + 2y = 11 8 - 8y + 2y = 11 8 - 6y = 11

OpenStudy (anonymous):

i don't understand the fraction part

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

oky wait let me check t

OpenStudy (anonymous):

okay :)

Nnesha (nnesha):

good so now subtract 8 both side

Nnesha (nnesha):